OpenStudy (anonymous):

4(x − 1) = 2(x + 2) + 6

OpenStudy (pottersheep):

7 :D

OpenStudy (pottersheep):

4x - 4 = 2x + 4 +6

OpenStudy (pottersheep):

put all variables on one side

OpenStudy (pottersheep):

4x - 2x = 4 + 6 + 4

OpenStudy (pottersheep):

2x = 14 x = 7!

OpenStudy (zeesbrat3):

4x-4=2x+4+6 4x-4=2x+10 -2x -2x 2x-4=10 +4 +4 2x=14 2/2 14/2 x=7

OpenStudy (pottersheep):

Do you understand?